36-303: Sampling, Surveys & Society HW03: Hand in in-class Thursday, Mar 25, 2010

\underline{Reminders:}

• Everyone has an approved II.6 (final survey proposal and IRB application) so you should go ahead and collect your data.

• Rough schedule of project due dates for remainder of term:

Meeting of Minds Poster submission deadline Apr 1.

III.7 Progress Report on Data Collection & Analysis (Apr 1, 6 & 8) [20-min Powerpoint presentation in class]

III.8 Rough Draft of Final Report (Due Apr 8) [a single pdf or doc file emailed to me]

“Final” Midterm Exam Tues Apr 13

III.9 Final presentations (Apr 22, 27 & 29); Final Report Due (Due Fri Apr 20) [a single pdf or doc file emailed to me]

• Reading:

– Review all of Groves Ch 4: Sample Design & Sampling Error

– Review Groves Ch 6: Nonresponse

– New: Groves Ch 10: Post-Survey Processing

– If you have Lohr, you could see some gory details in Ch 2 (SRS), Ch 4 (Stratified Sampling) and Ch 5 (Cluster Sampling)

Exercises to Turn In (there are 5 exercises):

1. Groves, Ch 4, pp. 134–135, #8

2. Groves, Ch 6, pp. 196–197, #2

3. Groves, Ch 6, p. 199, #9

4. Groves, Ch 10, pp. 342–343 #3. The weighted mean wanted in part (a) uses the “base weights” in the formula at the top of p. 324. Refer to sections 10.5.1 and 10.5.2 to answer the qualitative part of part (a).

5. Groves, Ch 10, p. 343, #6. Make sure you give at least 2 strengths and 2 weaknesses for each method (which might be the same or might be different between methods), and directly compare the strengths and weaknesses.